OpenStudy (sighn0more):

What are the steps I should use to solve this problem? 2 - 1/2 log(base)n 4 - log(base)n 5

OpenStudy (sighn0more):

Would I need to convert the 2 to a log first?

OpenStudy (anonymous):

\[2-\log_{n} 4^{\frac{ 1 }{ 2 }}-\log_{n} 5=2-\log_{n} \left( 2*5 \right)\]

OpenStudy (sighn0more):

Don't the rules of logarithms state that when logs are subtracted they must be divided?

OpenStudy (sighn0more):

Also, the back of my math textbook writes that the answer to this problem is 1/2. The problem is that I do not know how to get to that answer.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i do not see an equal sign there, so what exactly are you trying to "solve" maybe it means "combine as a single log"?

OpenStudy (sighn0more):

Hello! And it wants me to write the expression as either a single log or an integer

OpenStudy (misty1212):

ok

OpenStudy (misty1212):

that makes more sense

OpenStudy (misty1212):

we can start with \[\frac{1}{2}\log(4)=\log(\sqrt4)=\log(2)\]

OpenStudy (misty1212):

that since \[\frac{1}{2}\log(4)=\log(4^{\frac{1}{2}})\]

OpenStudy (sighn0more):

Okay, so we could re-write the problem as: 2 - logn2 - logn5 ?

OpenStudy (misty1212):

then you have \[2-\log(2)-\log(5)\]

OpenStudy (misty1212):

that is the same as \[2-(\log(2)+\log(5))\]

OpenStudy (misty1212):

so \[2-\log(10)\] after you multiply b

OpenStudy (misty1212):

can't thing of any further to go, since you do not know the base

OpenStudy (misty1212):

it is \[\log_n(10)\] right? without knowing \(n\) you cannot get a number out of this

OpenStudy (sighn0more):

Oh! I see. I forgot to distribute the negative sign. Hmmm.....yep! That seems to be right.

OpenStudy (misty1212):

ok good not \(\frac{1}{2}\) for sure, unless you are told what exactly \(n\) is equal to

OpenStudy (sighn0more):

Definitely! Sorry, that was an error on my part. I was looking at the wrong problem. Thank you!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]